# 2.5cm2.5cm2cm2.0cm

### v1

@@ 5 Experimental results from SPS through RHIC to LHC
After the above phenomenological considerations, let us discuss recent experimental results as well. These can be compared to the first thorough two-pion Lévy HBT analysis in heavy-ion collisions, the one detailed in Ref. [6]. In this paper $\sqrt{s_{{}_{NN}}}=200$ GeV Au+Au collisions with 0-30% centrality were analyzed, and $\alpha$ values around 1.2 were found, with a close to negligible dependence on $m_{T}$. These were later confirmed in a preliminary 3D analysis [42]. In both analyses [6, 42] a decreasing trend of the Lévy HBT scale $R$ was found with $m_{T}$, a feature usually attributed to the expansion of the fireball, as this leads to smaller sources for particles with larger momenta [46, 47, 48]. This, however, was calculated and predicted for Gaussian HBT radii. Hence, it is an open question why this feature of $R\propto 1/\sqrt{m_{T}}$ holds for Lévy HBT radii as well. In the mentioned analysis of Ref. [6] an interesting behavior of the correlation strength parameter $\lambda$ was found: it exhibits a “hole” (i.e., a strong decrease) in the small-$m_{T}$ region, attributed to an in-medium mass decrease of the $\eta^{\prime}$ meson (see Ref. [6] for details). These are the features of the Lévy HBT parameters ($\alpha$, $R$, $\lambda$) that shall be investigated in recent results as well.
The detailed centrality and average transverse mass dependence of these two-pion Lévy HBT parameters has been measured in STAR at $\sqrt{s_{{}_{NN}}}=200$ GeV, and reported recently by D. Kincses at the 52nd International Symposium on Multiparticle Dynamics and the XVI. Workshop of Particle Correlations and Femtoscopy. These results are summarized in the same special issue as this paper. We just note here that a similar $\alpha$ value was found as in Ref. [6], as shown in the summary plot of Fig. 4, and similar features for the $m_{T}$-dependence of $R$ and $\lambda$ parameters were found as in Ref. [6].
A similar analysis in the CMS experiment in $\sqrt{s_{{}_{NN}}}=5.02$ TeV Pb+Pb collisions was also performed (albeit without particle identification, under the assumption that most detected particles are pions), as reported in Ref. [49]. Here, it was found that the Lévy scale $R$ follows a similar trend versus $m_{T}$ as present in the RHIC results discussed above. It was also found that the exponent $\alpha$ is closer to the Gaussian case but clearly not equal to it. Furthermore, $\alpha$ values were found to be larger (closer to 2) for central collisions, this behavior is distinctly different from what was found in STAR. Fig. 4 also shows a central and a peripheral $\alpha$ measurement for both STAR and CMS, where the opposite trend is apparent.
@@
As mentioned, Fig. 4 compares the mean $\alpha$ values at several collision energies from SPS through RHIC to LHC. Interestingly, the 5.02 TeV Pb+Pb results and the 16.82 GeV Ar+Sc results yield the largest $\alpha$ values, while those for 200 GeV Au+Au and 16.82 GeV Be+Be collisions are significantly smaller. It is important to note however that the kinematic range of these experiments is also quite different: while there is essentially no low-momentum cut in NA61/SHINE (as in the fixed-target setup, small momentum particles can also be detected), there is one of about 0.2 GeV (in transverse momentum) in the RHIC experiments, while in case of CMS, particles with transverse momenta smaller than 0.5 GeV are discarded. This may also strongly influence the results of Fig. 4.
@@
[…]
@@ 6 Discussion and conclusions
We have detailed several phenomenological and experimental results related to femtoscopy with Lévy sources. In light of these, the natural question arises: is it advantageous to utilize Lévy distributed sources in describing momentum correlations? We have shown above that the appearing extra parameter, the Lévy-exponent $\alpha$ has a physical meaning for each of the possible physical reasons for its appearance (jet fragmentation, critical phenomena, anomalous diffusion, resonance decays). In order to interpret it, one needs to disentangle the above effects. Equally importantly, when measuring source size ($R$) or strength ($\lambda$) parameters with a Gaussian assumption, the actual size and shape information are entangled, and the more general Lévy assumption leads to refined results. One shall also note that the source, of course, does not necessarily exhibit an exact Lévy shape – deviations from it can be investigated via the method of Lévy expansion [54], and in Ref. [6] all expansion terms turned out to be zero, i.e., no deviation from the Lévy shape was found. This is not necessarily true, however, for all other experiments, and a similar expansion analysis can be performed in those cases as well.
@@
It is furthermore an important question whether utilizing the imaging technique mentioned above and detailed in Ref. [43] or the related technique of spherical decomposition [55] makes the Lévy source assumption superfluous. Indeed these methods help us map out the particle emitting source. In order to quantify the collision energy or centrality evolution of the source (for example, its size), one then would proceed to determine source measures such as HWHI, HWHM (or even a root-mean-square) – or fit the imaged sources with a suitable shape, such as a Gaussian or a Lévy distribution. If this shape is indeed describing the data, then it is equivalent to fit the source or the correlation function. However, in the case of the correlation function, the uncertainties are more clear; hence, even in that case, it is advantageous to fit this experimental object. Hence the utilization of the Lévy sources (possibly with the Lévy expansion technique) has no real disadvantage compared to imaging as well.
In summary, Lévy sources have been observed from SPS through RHIC to LHC. The possible reasons for the appearance of Lévy sources in high-energy physics include, depending on system size and collision energy, jet fragmentation, critical phenomena, resonance decays, and anomalous diffusion, which need to be disentangled. In measurements Lévy exponent values between 1 and 2 were found, with a nontrivial collision energy and system size dependence, suggesting that multiple phenomena contribute to the observed $\alpha$. The particle type dependence was also found to be nontrivial, not following the simplest predictions based on anomalous diffusion in the hadronic stage. The Lévy scale (or HBT radius) $R$ was found to exhibit an $m_{T}$-dependence that was predicted for Gaussian radii: a feature yet to be understood via model calculations. Finally, femtoscopy with Lévy sources provides a method to accurately measure the correlation strength $\lambda$, and differences were found in the low-$m_{T}$ behavior between RHIC and SPS, possibly hinting at an effect being absent at SPS energies. These results underline the importance of considering Lévy shapes when describing momentum correlation functions in high-energy physics and the need for more detailed phenomenological studies to explore all factors contributing to the shape of the particle emitting source.
@@ Acknowledgments
We acknowledge the support of NKFIH grants K-138136, TKP2021-NKTA-64, and PD-146589. We also thank M. Nagy for insightful discussions.

### v3

@@ 5 Experimental Results from SPS through RHIC to LHC
After the above phenomenological considerations, let us discuss recent experimental results as well. These can be compared to the first thorough two-pion Lévy HBT analysis in heavy-ion collisions, as detailed in Ref. [6]. In this paper, $\sqrt{s_{{}_{NN}}}=200$ GeV Au+Au collisions with 0–30% centrality were analyzed, and $\alpha$ values around 1.2 were found, with a close to negligible dependence on $m_{T}$. These were later confirmed in a preliminary 3D analysis [42]. In both analyses [6, 42], a decreasing trend of the Lévy HBT scale $R$ was found with $m_{T}$, a feature usually attributed to the expansion of the fireball, as this leads to smaller sources for particles with larger momenta [46, 47, 48]. This, however, was calculated and predicted for Gaussian HBT radii. Hence, it is an open question why this feature of $R\propto 1/\sqrt{m_{T}}$ holds for Lévy HBT radii as well. In the mentioned analysis of Ref. [6], an interesting behavior of the correlation strength parameter $\lambda$ was found: it exhibits a “hole” (i.e., a strong decrease) in the small-$m_{T}$ region, attributed to an in-medium mass decrease of the $\eta^{\prime}$ meson (see Ref. [6] for details). These are the features of the Lévy HBT parameters ($\alpha$, $R$, $\lambda$) that shall be investigated in recent results as well.
The detailed centrality and average transverse mass dependence of these two-pion Lévy HBT parameters have been measured in Au+Au collisions with STAR at $\sqrt{s_{{}_{NN}}}=200$ GeV and reported recently by D. Kincses at the 52nd International Symposium on Multiparticle Dynamics and the XVI. Workshop of Particle Correlations and Femtoscopy, and in Ref. [49]. These results are summarized in the same Special Issue as this paper. We note here that a similar $\alpha$ value was found as in Ref. [6], as shown in the summary plot of Figure 4, and similar features for the $m_{T}$-dependence of $R$ and $\lambda$ parameters were found as in Ref. [6].
A similar analysis in the CMS experiment in $\sqrt{s_{{}_{NN}}}=5.02$ TeV Pb+Pb collisions was also performed (albeit without particle identification, under the assumption that most detected particles are pions), as reported in Ref. [50]. Here, it was found that the Lévy scale $R$ follows a similar trend versus $m_{T}$ as present in the RHIC results discussed above. It was also found that the exponent $\alpha$ is closer to the Gaussian case but clearly not equal to it. Furthermore, $\alpha$ values were found to be larger (closer to 2) for central collisions; this behavior is distinctly different from what was found in STAR. Figure 4 also shows a central and a peripheral $\alpha$ measurement for both STAR and CMS, where the opposite trend is apparent.
@@
[…]
As mentioned, Figure 4 compares the mean $\alpha$ values at several collision energies from SPS through RHIC to LHC. Interestingly, the 5.02 TeV Pb+Pb results and the 16.82 GeV Ar+Sc results yield the largest $\alpha$ values, while those for 200 GeV Au+Au and 16.82 GeV Be+Be collisions are significantly smaller. It is important to note, however, that the kinematic range of these experiments is also quite different: While there is essentially no low-momentum cut in NA61/SHINE (as in the fixed-target setup, small momentum particles can also be detected), there is one of about 0.2 GeV (in transverse momentum) in the RHIC experiments, while in case of CMS, particles with transverse momenta smaller than 0.5 GeV are discarded. This may also strongly influence the results of Figure 4.
@@
The $R$ parameter obtained in the above-mentioned experimental analyses exhibits a typical and characteristic dependence on $m_{T}$: it decreases significantly as $m_{T}$ increases, due to the non-static nature of the source and the fact that $R$ is not a geometric size but a kind of measure of homogeneity, as mentioned in the introduction. Hence, $m_{T}$-averaging would not be meaningful. To facilitate a comparison, in Figure 5, we show $R$ values from SPS [51, 52] and RHIC [6, 49] at $m_{T}=0.3$ GeV/$c^{2}$, along with an extrapolation of the CMS data [50] to this $m_{T}$ value, using the formula for $R(m_{T})$ utilized in Ref. [49]. We furthermore compare RHIC and LHC results at $m_{T}=0.6$ GeV/$c^{2}$ (extrapolation of the SPS data to 0.6 GeV/$c^{2}$ is less feasible due to the larger uncertainties there). It is clear that $R$ decreases with $m_{T}$, as detailed in the above-cited experimental papers as well. It is also clear that the Lévy scale depends on the initial system size; in particular, $R$ for Be+Be is smaller than for Ar+Sc at the same energy. There is also an increase from SPS to RHIC, partly probably due to the system size change but also due to the increased energy. Rather surprisingly, however, Lévy scales at the LHC are not larger than those at RHIC but slightly smaller. This may be connected to the larger $\alpha$ measured at LHC and the anti-correlation of $R$ and $\alpha$, as demonstrated by Figure 7c of Ref. [6], Figure 5 of Ref. [36], and Figures 4 and 5 of Ref. [9].
Figure 5: Experimental results on the $R$ value (at a mean $m_{T}$ of 0.3 and 0.6 GeV/$c^{2}$), as a function of collision energy ($\sqrt{s_{{}_{NN}}}$). The source of the data is the same as for Figure 4. The CMS data point for $m_{T}=0.3$ GeV/$c^{2}$ was obtained via an extrapolation, as noted in the text.
The third important physical parameter of HBT correlation functions is the correlation strength parameter $\lambda$, as introduced in the first section. It was observed at RHIC [6] that there is a decrease of $\lambda(m_{T})$ at small $m_{T}$ values. The core–halo model interpretation utilized in Ref. [6] suggests that this may indirectly connect to an in-medium mass modification of the $\eta^{\prime}$ meson. It is, hence, quite interesting that this behavior is absent at SPS, as if this effect could be “turned off” when decreasing the collision energy from RHIC to SPS. Investigating this (the $m_{T}$-dependence of $\lambda$) at various energies also in data from the RHIC beam energy scan will be important and interesting. Note that no particle identification was employed in the CMS analysis of Ref. [50]. Hence, $\lambda$ cannot be interpreted in terms of the core–halo model only. In Ref. [50], an attempt was made to correct for the lack of particle identification, and a new parameter $\lambda^{*}$ was shown. However, the accessible $m_{T}$-range of CMS does not allow for investigating the phenomenon mentioned above (i.e., the decrease of $\lambda$ at small $m_{T}$).
@@ 6 Discussion and Conclusions
We have detailed several phenomenological and experimental results related to femtoscopy with Lévy sources. In light of these, the natural question arises: is it advantageous to utilize Lévy distributed sources in describing momentum correlations? We have shown above that the appearing extra parameter, the Lévy exponent $\alpha$ has a physical meaning for each of the possible physical reasons for its appearance (jet fragmentation, critical phenomena, anomalous diffusion, resonance decays). In order to interpret it, one needs to disentangle the above effects. Equally importantly, when measuring source size ($R$) or strength ($\lambda$) parameters with a Gaussian assumption, the actual size and shape information are entangled, and the more general Lévy assumption leads to refined results. One shall also note that the source, of course, does not necessarily exhibit an exact Lévy shape—deviations from it can be investigated via the method of Lévy expansion [55], and in Ref. [6] all expansion terms turned out to be zero, i.e., no deviation from the Lévy shape was found. This is not necessarily true, however, for all other experiments, and a similar expansion analysis can be performed in those cases as well.
@@
It is furthermore an important question whether utilizing the imaging technique mentioned above and detailed in Ref. [43] or the related technique of spherical decomposition [56] makes the Lévy source assumption superfluous. Indeed, these methods help us map out the particle emitting source. In order to quantify the collision energy or centrality evolution of the source (for example, its size), one then would proceed to determine source measures such as HWHI and HWHM (or even a root-mean-square)—or fit the imaged sources with a suitable shape, such as a Gaussian or a Lévy distribution. If this shape indeed describes the data, then it is equivalent to fitting the source or the correlation function. However, in the case of the correlation function, the uncertainties are more clear; hence, even in that case, it is advantageous to fit this experimental object. Therefore, the utilization of the Lévy sources (possibly with the Lévy expansion technique) has no real disadvantage compared to imaging as well.
In summary, Lévy sources have been observed from SPS through RHIC to LHC. The possible reasons for the appearance of Lévy sources in high-energy physics include, depending on system size and collision energy, jet fragmentation, critical phenomena, resonance decays, and anomalous diffusion, which need to be disentangled. In measurements, Lévy exponent values between 1 and 2 were found, with a nontrivial collision energy and system size dependence, suggesting that multiple phenomena contribute to the observed $\alpha$. The particle type dependence was also found to be nontrivial, not following the simplest predictions based on anomalous diffusion in the hadronic stage. The Lévy scale (or HBT radius) $R$ was found to exhibit an $m_{T}$-dependence that was predicted for Gaussian radii: a feature yet to be understood via model calculations. Finally, femtoscopy with Lévy sources provides a method to measure the correlation strength $\lambda$ accurately, and differences were found in the low-$m_{T}$ behavior between RHIC and SPS, possibly hinting at an effect being absent at SPS energies. These results underline the importance of considering Lévy shapes when describing momentum correlation functions in high-energy physics and the need for more detailed phenomenological studies to explore all factors contributing to the shape of the particle emitting source.
@@ Acknowledgments
We acknowledge the support of NKFIH grants K-138136, TKP2021-NKTA-64, K-146913, and PD-146589. We also thank M. Nagy for insightful discussions.